The operating expense budget of MountainTop Creations is comprised of the following budget formulas for its fixed and variable selling and administrative expense items: Selling expenses: $288,000 per year + $18 per unit Administrative expenses: $552,000 per year + 2% of sales If MountainTop’s sales budget includes sales of 5,000 units each month at a price $50 per unit, the total budgeted operating expense for the year would be:
Answer #1

Selling expenses

= 288,000 + (18*12*5000 units)

= 1,368,000

Administrative expenses

= 552,000 + (2% * 5000*12*50)

= 612,000

Total = 1,368,000 + 612,000

= 1,980,000